# CAST: Component-Aligned 3D Scene Reconstruction from an RGB Image

### v1

@@ Abstract
Recovering high-quality 3D scenes from a single RGB image is a challenging task in computer graphics. Current methods often struggle with domain-specific limitations or low-quality object generation. To address these, we propose CAST (Component-Aligned 3D Scene Reconstruction from a Single RGB Image), a novel method for 3D scene reconstruction and recovery. CAST starts by extracting object-level 2D segmentation and relative depth information from the input image, followed by using a GPT-based model to analyze inter-object spatial relationships. This enables the understanding of how objects relate to each other within the scene, ensuring more coherent reconstruction. CAST then employs an occlusion-aware large-scale 3D generation model to independently generate each object’s full geometry, using MAE and point cloud conditioning to mitigate the effects of occlusions and partial object information, ensuring accurate alignment with the source image’s geometry and texture. To align each object with the scene, the alignment generation model computes the necessary transformations, allowing the generated meshes to be accurately placed and integrated into the scene’s point cloud. Finally, CAST incorporates a physics-aware correction step that leverages a fine-grained relation graph to generate a constraint graph. This graph guides the optimization of object poses, ensuring physical consistency and spatial coherence. By utilizing Signed Distance Fields (SDF), the model effectively addresses issues such as occlusions, object penetration, and floating objects, ensuring that the generated scene accurately reflects real-world physical interactions. Experimental results demonstrate that CAST significantly improves the quality of single-image 3D scene reconstruction, offering enhanced realism and accuracy in scene recovery tasks. CAST has practical applications in virtual content creation, such as immersive game environments and film production, where real-world setups can be seamlessly integrated into virtual landscapes. Additionally, CAST can be leveraged in robotics, enabling efficient real-to-simulation workflows and providing realistic, scalable simulation environments for robotic systems.
@@
Figure 1: CAST brings diverse 3D scenes to life from a single image, where the relationships between objects shaped by their physical roles and interactions come together to form a cohesive and immersive virtual environment.
footnotetext: $\star$ Equal contribution. $\dagger$ Project Leader.
1 Introduction
@@
Humans exist within clear networks of relationships—family, friends, coworkers—that guide our decisions and behaviors. These connections shape our world and give it structure. Similarly, objects in a space also function within their own networks [41], but less noticed. They do not just exist in isolation; their placement, design, and material arise from physical constraints, functional roles, and human design intentions and influence how we move, interact, and perceive space. For example, a chair leans against a table for support, a cup rests on a saucer, and a lamp’s light interacts with surrounding surfaces, casting shadows that shape the overall scene. Recognizing these relationships is critical for accurate scene parsing, modeling, and, more recently, 3D generation, ensuring virtual environments feel as realistic and coherent as the real world.
@@
Significant progress has been made in generating single objects from text or image prompts. Neural rendering approaches [61, 75] optimize implicit representations, while native 3D generators [90, 92, 79] directly create 3D shapes and textures via end-to-end learning. While these methods show promise for individual objects, applying them to generate entire scenes by assembling objects sequentially results in notable shortcomings. A key challenge is accurate pose estimation. Existing methods often assume objects are view-aligned, which is rarely the case in real-world scenes. Objects may appear in diverse orientations, constrained by design, physics, or partial occlusion. Yet, most existing methods prioritize geometric fidelity over pose alignment, leaving this critical aspect underexplored.
@@
An even more fundamental issue arises from the lack of inter-object spatial relationships. Even with accurate poses, generated scenes often suffer from physically implausible artifacts: objects penetrate one another, float, or fail to make contact where necessary. These errors stem from the absence of spatial and physical constraints that naturally bind objects together, much as human relationships structure our social world. While some recent methods [45, 91] encode spatial relationships implicitly using encoder-decoder architectures, they remain limited to specific domains such as indoor scenes. Other scene-level generators [21] position objects in a global coordinate system but neglect their relative poses and dependencies, further compromising realism and usability for downstream applications like editing, animation, and simulation.
@@
To this end, we propose CAST, a Component-Aligned 3D Scene reconsTruction method for compositional reconstruction of a 3D scene from a single RGB image. CAST generates high-quality 3D meshes for individual objects, along with their similarity transformations (rotation, translation, scale), ensuring alignment with the reference image and enforcing physically sound interdependencies. CAST starts by processing an unstructured RGB image using 2D foundation models (e.g., Florence-2 [80], GroundingDINO [48], SAM [63], Grounded-SAM [64]) to recognize, localize, and segment objects in an open-vocabulary manner. Off-the-shelf monocular depth estimators [74] provide partial 3D point clouds and initial estimates of inter-object spatial relationships, including relative transformations and scales.
@@
The first core component of CAST is our perceptive 3D instance generator with two modules: an occlusion-aware object generation module and a pose alignment generation module. The object generation module employs a latent diffusion-based generative model to produce high-fidelity object meshes conditioned on partial image segments and optional point clouds. This module incorporates an occlusion-aware 2D image encoder capable of inferring occluded regions, ensuring robust feature extraction for image conditions. To improve robustness to real-world point cloud conditioning, we simulate partial point clouds with occluded regions during training, enabling the model to handle occlusion effectively. The pose alignment module features an alignment generative model that produces a transformed partial point cloud, aligning with the complete geometry implicitly represented in the latent space. The similarity transformation is derived from the generated transformed point cloud and the partial point cloud estimated from the camera. Unlike direct pose regression methods [34, 38], our method estimates transformations through generation, capturing the multi-modal nature of pose alignment.
@@
The second core component of CAST addresses inter-object spatial relationships. Despite accurate pixel alignment, physically implausible artifacts such as penetration or floating can occur without explicit modeling of physical constraints. CAST introduces a physics-aware correction process to ensure spatial and physical coherence. GPT-4v [1] is utilized to identifies commonsense physical relationships grounded in the input image, which are then used to optimize object poses based on these constraints. This process ensures that reconstructed scenes exhibit realistic physical interdependencies, making them suitable for applications like simulation, editing, and rendering.
@@
Remarkably, CAST excels at generating perceptually realistic 3D scenes from a wide range of images, whether they are sourced from indoor or outdoor settings, real-world captured, or AI-generated. Unlike previous approaches [45, 18], CAST supports open-vocabulary reconstruction, even for challenging, in-the-wild images, thanks to our deliberate pipeline design. Quantitatively, CAST surpasses strong baselines in the indoor dataset, 3D-Front [22], regarding object- and scene-level geometry quality. It also outperforms on perceptual and physical realism across a diverse set of images, including in-the-wild scenarios, as verified by visual-language models and user studies.
@@
Given only a single image, CAST can faithfully reconstruct the scene, with detailed geometry, vivid textures of the objects, and more importantly, the spatial and physical interdependencies between them. This capability democratizes virtual creation: a single snapshot of a room or outdoor space becomes a fully realized 3D environment, where objects are precisely posed, interact naturally, and account for occlusions. Game developers can integrate real-world setups into immersive landscapes, and filmmakers can effortlessly generate intricate virtual sets—unlocking creative potential. Beyond entertainment, CAST paves the way for smarter robots. It can facilitate the real-to-simulation pipeline [43, 71] by enabling robotics researchers to construct digital replicas from real-world demonstration datasets with more efficient and scalable simulation workflows.
2 Related Work
@@
Photogrammetry is a widely used method to capture the physical world in high detail and translate it into digital form [27, 11, 56, 3, 4, 57, 35], but it requires tens to hundreds of images from multiple viewpoints, making it time-consuming, resource-intensive, and hard to scale. In contrast, single-image-based approaches are more efficient and scalable, requiring only one image that can be easily obtained from online repositories, eliminating the need for expensive scanning devices or multi-view setups.
2.1 Single Image Scene Reconstruction
@@
Scene-level reconstruction from a single image presents challenges due to object diversity, occlusions, and the need to preserve spatial relationships. A starting point is monocular depth estimation, where depth is inferred from a single image, typically generating a depth point cloud [86, 7, 60, 74, 84]. While it provides valuable information, it struggles with occlusions and hidden portions of the scene. To address this, novel view synthesis methods use representations like radiance fields [87, 70, 89] and 3D Gaussian [67, 66], learning occlusion priors from 3D datasets [25, 17, 10, 65]. Despite these advances, monocular reconstruction methods still often struggle to provide detailed and precise scene representations.
@@
Some methods, such as 3DFront [22], focus on directly regressing geometries along with their semantic labels in the scene [16, 26, 15, 12]. These approaches typically rely on scene datasets with ground truth object annotations, such as Matterport3D [22] and 3DFront [22], which are often small in scale and limited to indoor room environments. However, the feed-forward nature of these methods leads to the generation of geometries that often lack sufficient detail and quality.
@@
To better film real world to digital, other methods turn to retrieval-based approaches [40, 28, 37, 23, 18], which enhance scene quality by searching for and replacing objects in a scene with similar objects from a pre-existing dataset. These methods incorporate advanced tools such as GPT-4 [1], SAM [36, 64], and depth priors to decompose scenes. While these methods improve scene realism by integrating real-world objects, they are constrained by the richness and scope of the datasets they rely on. For scenes outside the dataset’s domain, retrieval-based methods either produce erroneous results or fail to find suitable replacements, significantly degrading the quality of the reconstructed scene.
@@
Figure 2: Overview of the proposed pipeline. The input RGB image is processed through scene analysis to extract key information, followed by pose-aware generation to create initial 3D models. Physical constraint refinement ensures realistic interactions and spatial relationships, yielding a high-quality, mesh-based 3D scene.
2.2 Reconstruction as Generation
@@
Much of the current research in 3D asset generation focuses on distilling 3D geometry from 2D images generative models  [61, 69, 75]. More recent developments expand this approach by incorporating multi-view images for supervision [50, 47, 46, 73, 51, 77], often trained on rendered images from large-scale object datasets like Objaverse [20], to enhance view consistency during generation. Some approaches directly regress the shape and appearance of individual objects based on input image [32, 68]. While these methods achieve satisfactory visual results, they frequently fail to reproduce fine geometric details. To improve the quality of 3D geometry, a growing body of work has moved away from 2D supervision entirely, opting instead to train directly on 3D assets [20, 19]. These methods produce high-quality object-level geometries with advanced processing techniques [90, 92, 79]. However, such approaches focus on isolated objects and fail to address scene-level challenges, such as modeling spatial hierarchies, inter-object relationships, and environment lighting. Scene generation remains underdeveloped due to the high computational and representational complexity of modeling object relationships, lighting, and materials. Despite progress, current approaches still struggle to produce fully realized, editable 3D scenes. Existing paradigms either use video diffusion models [31, 30, 8] to generate navigable 2D projections [9, 88], or rely on diffusion priors for volumetric scene approximations via 3D Gaussian splatting [78, 24, 44]. While these methods yield compelling visuals, they are incompatible with traditional production pipelines, lacking editable meshes, UV mappings, and decomposable PBR materials.
@@
A more feasible paradigm decomposes scenes into modular components—objects, backgrounds, and environmental generating and reassembling them into an editable scene graph for greater flexibility and precision. For example, Gen3DSR [21] uses DreamGaussian [69] for open-vocabulary reconstruction. However, it struggles with occlusions, pose estimation, and editing individual objects, while relying on 2D models leads to poor geometric details and low-fidelity representations. Another recent work, Midi [33], learns spatial relationships between objects in a scene but requires training on datasets with ground truth 3D meshes and annotations. This reliance on specific datasets limits its scalability and generalization to arbitrary scenes. In this paper, we present a novel scene reconstruction pipeline that generates each object independently and aligns them to form a cohesive scene. Unlike existing methods, our approach preserves accurate geometry, textures, and consistent spatial relationships, resulting in more realistic, reliable, and editable reconstructions, with improved quality and flexibility. Our approach generates each object independently and aligns them into a cohesive scene, preserving accurate geometry, textures, and consistent spatial relationships, resulting in more realistic, reliable, and editable reconstructions with improved quality and flexibility.
2.3 Physics-Aware 3D Modeling

### v2

@@ Abstract.
Recovering high-quality 3D scenes from a single RGB image is a challenging task in computer graphics. Current methods often struggle with domain-specific limitations or low-quality object generation. To address these, we propose CAST (Component-Aligned 3D Scene Reconstruction from a Single RGB Image), a novel method for 3D scene reconstruction. CAST starts by extracting object-level 2D segmentation and relative depth information from the input image, followed by using a GPT-based model to analyze inter-object spatial relations. This enables understanding of how objects relate to each other within the scene, ensuring more coherent reconstruction. CAST then employs an occlusion-aware large-scale 3D generation model to independently generate each object’s full geometry, using Masked Auto Encoder (MAE) and point cloud conditioning to mitigate the effects of occlusions and partial object information, ensuring accurate alignment with the source image’s geometry and texture. To align each object with the scene, the alignment generation model computes the necessary transformations, allowing the generated meshes to be accurately placed and integrated into the scene’s point cloud. Finally, CAST applies a physics-aware correction mechanism, which leverages a fine-grained relation graph to generate a constraint graph. This graph guides the optimization of object poses, ensuring physical consistency and spatial coherence. By utilizing Signed Distance Fields (SDF), the model effectively addresses issues such as occlusions, object penetration, and floating objects, ensuring that the generated scene accurately reflects real-world physical interactions. Experimental results demonstrate that CAST significantly improves the quality of single-image 3D scene reconstruction, offering enhanced realism and accuracy in scene understanding and reconstruction tasks. CAST has practical applications in virtual content creation, such as immersive game environments and film production, where real-world setups can be seamlessly integrated into virtual landscapes. Additionally, CAST can be leveraged in robotics, enabling efficient real-to-simulation workflows and providing realistic, scalable simulation environments for robotic systems.
Open-Vocabulary Scene Reconstruction, Generative Pose Alignment, Occlusion-aware 3D Generation, Physical Consistency
submissionid: 364ccs: Computing methodologies Artificial intelligence
@@
Figure 1. CAST brings diverse 3D scenes to life from a single image, where the relations between objects shaped by their physical roles and interactions come together to form a cohesive and immersive virtual environment.
1. Introduction
@@
Humans exist within clear networks of relations—family, friends, coworkers—that guide our decisions and behaviors. These connections shape our world and give it structure. Similarly, objects in a space also function within their own networks (Latour, 2005), but less noticed. They do not just exist in isolation; their placement, design, and material arise from physical constraints, functional roles, and human design intentions and influence how we move, interact, and perceive space. For example, a chair leans against a table for support, a cup rests on a saucer, and a lamp’s light interacts with surrounding surfaces, casting shadows that shape the overall scene. Recognizing these relations is critical for accurate scene parsing, modeling, and, more recently, 3D generation, ensuring virtual environments feel as realistic and coherent as the real world.
@@
Significant progress has been made in generating single objects from text or image prompts. Neural rendering approaches (Poole et al., 2022; Wang et al., 2024a) optimize implicit representations, while native 3D generators (Zhang et al., 2023, 2024a; Xiang et al., 2024) directly create 3D shapes and textures via end-to-end learning. While these methods show promise for individual objects, applying them to generate entire scenes by assembling objects sequentially faces with notable shortcomings. A key challenge is accurate pose estimation. Existing methods often assume objects are view-aligned, which is rarely the case in real-world scenes. Objects may appear in diverse orientations, constrained by design, physics, or partial occlusion. Yet, most existing methods prioritize geometric fidelity over pose alignment, leaving this critical aspect underexplored.
@@
An even more fundamental issue arises from the lack of inter-object spatial relations. Even with accurate poses, generated scenes often suffer from physically implausible artifacts: objects penetrate one another, float, or fail to make contact where necessary. These errors stem from the absence of spatial and physical constraints that naturally bind objects together, much as human relations structure our social world. While some recent methods (Liu et al., 2022; Zhang et al., 2024b) encode spatial relations implicitly using encoder-decoder architectures, they remain limited to specific domains such as indoor scenes. Other scene-level generators (Dogaru et al., 2024) position objects in a global coordinate system but neglect their relative poses and dependencies, further compromising realism and usability for downstream applications like editing, animation, and simulation.
@@
To this end, we propose CAST, a Component-Aligned 3D Scene reconstruction method for compositional reconstruction of a 3D scene from a single RGB image. CAST generates high-quality 3D meshes for individual objects, along with their similarity transformations (rotation, translation, scale), ensuring alignment with the reference image and enforcing physically sound interdependencies. CAST starts by processing an unstructured RGB image using 2D foundation models (e.g., Florence-2 (Xiao et al., 2024), GroundingDINO (Liu et al., 2025), SAM (Ravi et al., 2024), Grounded-SAM (Ren et al., 2024)) to recognize, localize, and segment objects in an open-vocabulary manner. Off-the-shelf monocular depth estimators (Wang et al., 2024b) provide partial 3D point clouds and initial estimates of inter-object spatial relations, including relative transformations and scales.
@@
The first core component of CAST is our perceptive 3D instance generator with two modules: an occlusion-aware object generation module and a pose alignment generation module. The object generation module employs a latent diffusion-based generative model to produce high-fidelity object meshes conditioned on partial image segments and optional point clouds. This module incorporates an occlusion-aware 2D image encoder capable of inferring occluded regions, ensuring robust feature extraction for image conditions. To improve robustness to real-world point cloud conditioning, we simulate partial point clouds with occluded regions during training, enabling the model to handle occlusion effectively. The pose alignment module features an alignment generative model that produces a transformed partial point cloud, aligning with the complete geometry implicitly represented in the latent space. The similarity transformation is derived from the generated transformed point cloud and the partial point cloud estimated from the camera. Unlike direct pose regression methods (Kehl et al., 2017; Labbé et al., 2020), our method estimates transformations through generation, capturing the multi-modal nature of pose alignment.
@@
The second core component of CAST addresses inter-object spatial relations. Despite accurate pixel alignment, physically implausible artifacts such as penetration or floating can occur without explicit modeling of physical constraints. CAST introduces a physics-aware correction process to ensure spatial and physical coherence. GPT-4v (Achiam et al., 2023) is utilized to identifies commonsense physical relations grounded in the input image, which are then used to optimize object poses based on these constraints. This process ensures that reconstructed scenes exhibit realistic physical interdependencies, making them suitable for applications like simulation, editing, and rendering.
@@
Remarkably, CAST excels at generating perceptually realistic 3D scenes from a wide range of images, whether they are sourced from indoor or outdoor settings, real-world captured, or AI-generated. Unlike previous approaches (Liu et al., 2022; Dai et al., 2024), CAST supports open-vocabulary reconstruction, even for challenging, in-the-wild images, thanks to our deliberate pipeline design. Quantitatively, CAST surpasses strong baselines in the indoor dataset, 3D-Front (Fu et al., 2021), regarding object- and scene-level geometry quality. It also outperforms on perceptual and physical realism across a diverse set of images, including in-the-wild scenarios, as verified by visual-language models and user studies.
@@
Given only a single image, CAST can faithfully reconstruct the scene, with detailed geometry, vivid textures of the objects, and more importantly, the spatial and physical interdependencies between them. This capability democratizes virtual creation: a single snapshot of a room or outdoor space becomes a fully realized 3D environment, where objects are precisely posed, interact naturally, and account for occlusions. Game developers can integrate real-world setups into immersive landscapes, and filmmakers can effortlessly generate intricate virtual sets—unlocking creative potentials. Beyond entertainment, CAST paves the way for smarter robots. It can facilitate the real-to-simulation pipeline (Li et al., 2024a; Torne et al., 2024) by enabling robotics researchers to construct digital replicas from real-world demonstration datasets with more efficient and scalable simulation workflows.
2. Related Work
@@
Photogrammetry is a widely used method to capture the physical world in high detail and translate it into digital form (Goesele et al., 2007; Chen et al., 2018; Mildenhall et al., 2020; Barron et al., 2021, 2022; Müller et al., 2022; Kerbl et al., 2023), but it requires tens to hundreds of images from multiple viewpoints, making it time-consuming, resource-intensive, and hard to scale. In contrast, single-image-based approaches are more efficient and scalable, requiring only one image that can be easily obtained from online repositories, eliminating the need for expensive scanning devices or multi-view setups.
2.1. Single Image Scene Reconstruction
@@
Scene-level reconstruction from a single image presents challenges due to object diversity, occlusions, and the need to preserve spatial relations. A starting point is monocular depth estimation, where depth is inferred from a single image, typically generating a depth point cloud (Yin et al., 2023; Bhat et al., 2023; Piccinelli et al., 2024; Wang et al., 2024b; Yang et al., 2024b). While it provides valuable information, it struggles with occlusions and hidden portions of the scene. To address this, novel view synthesis methods use representations like radiance fields (Yu et al., 2021; Tian et al., 2023; Yu et al., 2022) and 3D Gaussian (Szymanowicz et al., 2024b, a), learning occlusion priors from 3D datasets (Geiger et al., 2013; Dai et al., 2017; Chang et al., 2015; Sun et al., 2018). Despite these advances, monocular reconstruction methods still often struggle to provide detailed and precise scene representations.
@@
Some methods focus on directly regressing geometries along with their semantic labels in the scene (Dahnert et al., 2021; Gkioxari et al., 2022; Chu et al., 2023; Chen et al., 2024a). These approaches typically rely on scene datasets with ground truth object annotations, such as Matterport3D (Fu et al., 2021) and 3DFront (Fu et al., 2021), which are often small in scale and limited to indoor room environments. However, the feed-forward nature of these methods leads to the generation of geometries that often lack sufficient detail and quality.
@@
To better film real world to digital, other methods turn to retrieval-based approaches (Langer et al., 2022; Gümeli et al., 2022; Kuo et al., 2021; Gao et al., 2024b; Dai et al., 2024), which enhance scene quality by searching for and replacing objects in a scene with similar objects from a pre-existing dataset. These methods incorporate advanced tools such as GPT-4 (Achiam et al., 2023), SAM (Kirillov et al., 2023; Ren et al., 2024), and depth priors to decompose scenes. While these methods improve scene realism by integrating real-world objects, they are constrained by the richness and scope of the datasets they rely on. For scenes outside the dataset’s domain, retrieval-based methods either produce erroneous results or fail to find suitable replacements, significantly degrading the quality of the reconstructed scene.
@@
Figure 2. Overview of the proposed pipeline. The input RGB image is processed through scene analysis to extract key information, followed by pose-aware generation to create initial 3D models. Physical constraint refinement ensures realistic interactions and spatial relations, yielding a high-quality, mesh-based 3D scene.
2.2. Reconstruction as Generation
@@
Much of the current research in 3D asset generation focuses on distilling 3D geometry from 2D images generative models  (Poole et al., 2022; Tang et al., 2023; Wang et al., 2024a). More recent developments expand this approach by incorporating multi-view images for supervision (Liu et al., 2023a, c, 2024; Voleti et al., 2025; Long et al., 2024; Wu et al., 2024a), often trained on rendered images from large-scale object datasets like Objaverse (Deitke et al., 2023), to enhance view consistency during generation. Some approaches directly regress the shape and appearance of individual objects based on input image (Hong et al., 2023; Tang et al., 2025). While these methods achieve satisfactory visual results, they frequently fail to reproduce fine geometric details. To improve the quality of 3D geometry, a growing body of work has moved away from 2D supervision entirely, opting instead to train directly on 3D assets (Deitke et al., 2023, 2024). These methods produce high-quality object-level geometries with advanced processing techniques (Zhang et al., 2023, 2024a; Xiang et al., 2024). However, such approaches focus on isolated objects and fail to address scene-level challenges, such as modeling spatial hierarchies, inter-object relations, and environment lighting. Scene generation remains underdeveloped due to the high computational and representational complexity of modeling object relations, lighting, and materials. Despite progress, current approaches still struggle to produce fully realized, editable 3D scenes. Existing paradigms either use video diffusion models (Ho et al., 2022b, a; Blattmann et al., 2023) to generate navigable 2D projections (Bruce et al., 2024; Yu et al., 2024), or rely on diffusion priors for volumetric scene approximations via 3D Gaussian splatting (Wu et al., 2024b; Gao et al., 2024a; Liang et al., 2024). While these methods yield compelling visuals, they are incompatible with traditional production pipelines, lacking editable meshes, UV mappings, and decomposable PBR materials.
@@
A more feasible paradigm decomposes scenes into modular components—objects, backgrounds, and environmental generating and reassembling them into an editable scene graph for greater flexibility and precision. For example, Gen3DSR (Dogaru et al., 2024) uses DreamGaussian (Tang et al., 2023) for open-vocabulary reconstruction. However, it struggles with occlusions, pose estimation, and editing individual objects, while relying on 2D models leads to poor geometric details and low-fidelity representations. Another recent work, Midi (Huang et al., 2024), learns spatial relations between objects in a scene but requires training on datasets with ground truth 3D meshes and annotations. This reliance on specific datasets limits its scalability and generalization to arbitrary scenes.
Our approach shares conceptual foundations with classical analysis-by-synthesis methods (Yuille and Kersten, 2006), as both aim to infer 3D structure by generating explanations for observed imagery. However, while analysis-by-synthesis relies on iterative rendering and pixel-level optimization, our method leverages pre-trained generative models and learned priors to synthesize plausible 3D scenes directly, often bypassing explicit rendering and optimization loops, thereby improving scalability, efficiency, and adaptability to open-world scenarios.
Building on this foundation, we present a novel scene reconstruction pipeline that generates each object independently and aligns them into a cohesive scene. Unlike existing methods, our approach preserves accurate geometry, textures, and consistent spatial relations, resulting in more realistic, reliable, and editable reconstructions with improved quality and flexibility.
2.3. Physics-Aware 3D Modeling